# HandS3C: 3D Hand Mesh Reconstruction with State Space Spatial Channel Attention from RGB images

### v3

@@ Abstract
Reconstructing the hand mesh from one single RGB image is a challenging task because hands are often occluded by other objects. Most previous works attempt to explore more additional information and adopt attention mechanisms for improving 3D reconstruction performance, while it would increase computational complexity simultaneously. To achieve a performance-reserving architecture with high computational efficiency, in this work, we propose a simple but effective 3D hand mesh reconstruction network (i.e., HandS3C), which is the first time to incorporate state space model into the task of hand mesh reconstruction. In the network, we design a novel state-space spatial-channel attention module that extends the effective receptive field, extracts hand features in the spatial dimension, and enhances regional features of hands in the channel dimension. This helps to reconstruct a complete and detailed hand mesh. Extensive experiments conducted on well-known datasets facing heavy occlusions (such as FREIHAND, DEXYCB, and HO3D) demonstrate that our proposed HandS3C achieves state-of-the-art performance while maintaining a minimal parameters.
1 Introduction
@@
The reconstruction of 3D hand mesh from one single RGB image has a wide range of applications in many fields, such as VR/AR, robotics, and human-computer interaction. 3D hand mesh reconstruction based on RGB images for the reconstruction has attracted much attention due to its low cost and computational friendliness. Influenced by the rapid development of deep learning, many excellent works Hasson2019LearningJR; Boukhayma20193DHS; Choutas2020MonocularEB; Hampali2021KeypointTS; Wang_2023; park2022handoccnet; li2022interacting have emerged in this field. These works focus on important issues related to hand-object occlusion and the loss of local information. Despite significant advancements, most methods still struggle to reconstruct hand poses in scenarios with substantial occlusion or when the prior knowledge is not sufficient.
@@
The previous methods can be broadly categorized into two categories. One is to incorporate additional prior knowledge. To cite few, Zimm et al. zimmermann2017learning obtained hand region features via a hand segmentation network. Wang et al. Wang_2023 added keypoint coordinates as strong prior information to constrain hand pose. These methods can enhance pose reconstruction performance, but lack essential prior information in practical applications. Another category is to learn the intrinsic connection between image features through deep networks, Park et al. park2022handoccnet extracted the complete hand information by distinguishing the features. Jiao et al. https://doi.org/10.1049/ell2.13040 used the Transformer architecture to obtain the global information and expect to learn the hand feature mapping from the occluded part. These methods tend to mine the association between image features, but there are still some limitations. First, the attention mechanism leads to excessive computational complexity. Second, an overload of global information can interfere with the effective receptive field luo2017understanding, as shown in Figure. 1 (b).
Based on the above observations, we propose a simple but effective 3D hand mesh reconstruction network called as HandS3C. This framework incorporates the state space model liu2024vmamba; ma2024umamba; ruan2024vmunet into the field of hand mesh reconstruction for the first time. At the same time, combining the idea of spatial-channel feature fusion hu2019squeezeandexcitation, State Space Spatial Channel (S3C) attention module is constructed to emphasize the contribution of channel-dimensional features. This not only increases the range of effective receptive fields in the network, but also effectively avoids the square-step computational complexity of traditional attention. Specifically, our main contributions are summarized as follows:
(1) We design a 3D hand mesh reconstruction network called as HandS3C, which introduces the state space model to the task of hand mesh reconstruction for the first time. This network can effectively improve hand reconstruction performance without the need for additional prior knowledge.
(2) We propose a spatial and channel-wise parallel scanning approach to compensate for the missing channel information in planar scanning. The state-space spatial-channel attention module is constructed, which can enhance the effective receptive field range while maintaining a small number of parameters.
(3) Our method achieves state-of-the-art on three datasets, FREIHAND, DEXYCB and HO3D, with a small number of parameters.
Figure 1: Effective receptive fields (from left to right: FPN, Self-Attention, State Space Attention (SSA) and State-Space Spatial-Channel Attention (S3C)).
2 Related work
2.1 Additional Information based Methods
It is difficult to restore 3D hand gestures from RGB image features alone, and most researchers use heatmaps Boukhayma20193DHS; lv2021handtailor; iqbal2018hand, segmentation zimmermann2017learning; zhang2021hand, depth maps rezaei2022trihornnet; rezaei2023pushing; 10.1007/978-3-030-01270-0_15, and density maps zhao2023semisupervised; chen2021i2uvhandnet to incorporate additional information to help the network better learn hand features in occlusion situations. In recent years, some scholars Wang_2023; jiang2023a2jtransformer incorporate human prior knowledge to guide the network to focus on the target region, Wang et al. Wang_2023 proposed HandGCAT, which learns the prior knowledge of the hand joint positions through the knowledge-guided graph convolution module, and then through the cross-attention Transformer to guide the network to inject joint information into the occluded region. In addition, other scholars Hasson2019LearningJR; wang2022interacting; qi2024hoisdf; li2022artiboost; chen2022tracking have argued that the object pose affects the hand pose. To end this, Qi et al. qi2024hoisdf introduced Signed Distance Field (SDF) to reconstruct the occluded object and the hand, and this method is able to determine the grip state of the hand by estimating the object pose and then determining the hand’s grasping state.
2.2 Attention based Methods
Attention mechanisms have achieved great success in the field of deep learning, due to their ability to map long-term contextual relevance. In the task of hand mesh reconstruction, occluded hand features can be learned through global mapping with various attention modules, which in turn can effectively solve the occlusion problem park2022handoccnet; li2022interacting; cho2022crossattention; feng2021collaborative; pavlakos2023reconstructing. In establishing the global context of hand-object interaction, Pavlakos et al. pavlakos2023reconstructing learn the mapping relationship between hand and occlusion from large-scale data through the Vision Transformer (ViT) architecture. Cho et al. cho2022crossattention obtained the features related to the region of interest from the occluded portion through cross-attention. Park et al. park2022handoccnet made full use of the information of occluded region as an enhancement of image feature of the hand and refine it to get the complete hand features.
2.3 State Space Model
With the proposal of attention mechanism vaswani2023attention, it makes more models can effectively focus on global information mapping, but its quadratic computational complexity leads to computational inefficiency in the case of long sequences. In the field of natural language processing, Gu et al. gu2022efficiently proposed the S4 model to optimize the state space model (SSM) in order to solve the problems of computational efficiency and long sequence dependency. However, the S4 model would process all inputs in the exactly same way, which is not in line with the practical intuition that each sequence should receive different attention. Therefore, Gu et al. gu2023mamba introduced a selection mechanism to the Mamba architecture (S6) for learning the parameters in the traditional state-space model, through which it is able to focus on or ignore some sequences. Compared to traditional attention computation methods, S6 enables each element in a 1-D array (e.g., a text sequence) to interact with any previously scanned samples through a compressed hidden state, effectively reducing the quadratic complexity to a linear one. Influenced by this idea, Liu et al. liu2024vmamba introduced it to vision tasks by combining the ViT model dosovitskiy2021image to transform images into sequences, and proposed a cross-scanning module to solve the problem of image orientation sensitivity. This is made possible by the selectivity of the state space model, which allows it to enhance similar features in the hidden space. Therefore, we abstract the selective state-space model and the cross-scanning module therein into a state space attention (SSA) module, as shown in Figure. 1 (c). In addition, some researchers ruan2024vmunet; ma2024umamba; Deng2024CUMambaSS have also combined state-space modeling with traditional CNN architectures to extract multi-scale global features in U-shaped structures by state-space selection modeling, and retained the deep-level information. In this paper, we introduce state space model into the field of hand mesh reconstruction for the first time , and design a state space spatial channel attention module (S3C). The module performs parallel scanning in both spatial and channel dimensions, and uses a selective state-space model to achieve feature selection in the hidden space. This method is able to extend the effective receptive field with a small number of parameters and focuses the network attention on the hand region, thus improving the accuracy of hand mesh reconstruction.
@@ 2.3 State Space Model
Figure 2: Overall architecture of HandS3C including Backbone, S3C and Regressor. HandS3C uses Resnet50 to extract image features and fuses multiscale features by upsampling layer by layer. It then uses S3C to obtain hand region features through parallel spatial and channel scanning. Finally, the regressor obtains keypoint information by heatmap encoding and outputs MANO parameters. The final 3D hand mesh is obtained by forwarding the MANO parameters to the MANO layer.
3 Methodology
3.1 Overview
@@
The overall network architecture is shown in Fig. 2 and consists of Backbone, S3C and Regressor components. Given a hand image $I\in\mathbb{R}^{256\times 256\times 3}$. First, the appearance information is extracted by ResNet-50 based encoder-decoder and the multi-scale features are fused layer by layer to get the original feature map $F\in\mathbb{R}^{256\times 32\times 32}$. Then, it is fed into the S3C module and scanned in both spatial and channel dimensions to compute a 2D feature map and fused to obtain enhanced hand features $H\in\mathbb{R}^{256\times 32\times 32}$. Finally, the hand features $H$ are passed to the Regressor to calculate the keypoint heat map $HM\in\mathbb{R}^{21\times 32\times 32}$ to predict the keypoint coordinates. For the final 3D hand mesh, we used the most popular MANO parametric hand MANO:SIGGRAPHASIA:2017. It is an articulated hand network generated from differentiable functions that takes as input the parameters that control the shape $\beta\in\mathbb{R}^{10}$ and pose $\theta\in\mathbb{R}^{48}$ of the hand.
3.2 Backbone
@@
In the original feature extraction part, we use ResNet-50 he2015deep to extract image information and fuse multi-scale features layer by layer lin2017feature in an U-shape encoder-decoder structure. Specifically, when the input image size is $256\times 256\times 3$, the size of the first layer of features after ResNet-50 becomes $128\times 128\times 128$. In the subsequent layers, each layer of image features $F_{i}\in\frac{H_{i}}{2^{i}}\times\frac{W_{i}}{2^{i}}\times C_{i},i\in(1,2,3,4),C_{i}\in(256,512,1024,2048)$ is obtained by downsampling. Then, upsampling is performed and the image features $F_{i}$ are incorporated layer by layer to get the last layer of feature extraction with size of $256\times 64\times 64$. Finally, the feature mapping map $F\in\mathbb{R}^{256\times 32\times 32}$ is obtained by fusing the feature information through average pooling.
3.3 State Space Spatial Channel Attention Module
State Space Models.
S4ND nguyen2022s4nd was the first study to introduce state space models into the field of vision by enabling it to receive 2D image information through a simple extension of the S4 model. The S6 module proposed by Gu et al. gu2023mamba was shown in VMamba liu2024vmamba to be able to capture the global feature information of 2D image features through a selective scanning mechanism enabling it to capture the global feature information of the 2D image features and maintain linear complexity. The state space model expects to map the input $x(t)\in\mathbb{R}^{L}$ to the output $y(t)\in\mathbb{R}^{L}$ through the implicit state $h(t)$ and the parameter $A\in\mathbb{R}^{N\times N},B,C\in\mathbb{R}^{N},D\in\mathbb{R}^{d}$ with the following equation:
@@ 3.3 State Space Spatial Channel Attention Module
In order to computationally achieve efficient operations gu2021combining, the ordinary differential equations above need to be transformed into discrete functions. Specifically, when the input size is $X_{i}\in\mathbb{R}^{\mathrm{L\times d}}$, the discretization is performed following the following rule:
$h_{i}=\overline{A}h_{i-1}+\overline{B}x_{i},\\ y_{i}=\overline{C}h_{i}+\overline{D}x_{i},\\ \overline{A}=e^{\Delta A},\\ \overline{B}=(e^{\Delta A}-I)A^{-1}B,\\ \overline{C}=C\$ (2)
where $\mathrm{B,C\in\mathbb{R}^{d\times N}}$ and $\Delta\in\mathbb{R}^{\mathrm{d}}$. In practice, we improve the $\bar{B}$ approximation by means of a first-order Taylor series gu2023mamba:
@@
Unlike traditional spatial state models, the S6 block introduces a selective scanning mechanism, where $B,C\in\mathbb{R}^{bs\times L\times N}$, and $\Delta\in\mathbb{R}^{bs\times L\times d}$ can be obtained from inputs of size $X\in\mathbb{R}^{bs\times L\times d}$. This means that S6 can sense contextual information from the inputs and update the weights.
Spatial-wise Selective Scanning.
Unlike textual information, image information contains 2D spatial information such as local texture and global structure, and it is difficult to obtain the contextual relationship of the local sense field of an image by scanning with a simple S6 module. The 1-D convolution is extended to 2-D convolution in S4ND gu2021combining, and this approach leads to the loss of the ability to dynamically update weights based on context. In VMamba liu2024vmamba, rows and columns are unfolded and scanned along in four different directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right, by means of a cross-scanning module. This allows any pixel to contain contextual information from all four directions and can enhance the size of the sensory field in the $H\times W$ plane. Sequences in each of the four directions are computed in the Spatial-wise Selective Scanning to obtain $B,C\in\mathbb{R}^{bs\times N\times L_{H\times W}}$, and $\Delta\in\mathbb{R}^{bs\times d\times L_{H\times W}}$, which in turn computes the feature map. Finally, the four sequences are merged to obtain a new sequence with the same size as the input sequence.
Channel-wise Selective Scanning.
Although spatial features such as local texture and global structure intrinsic to 2D image features are already available through Liu et al liu2024vmamba. However, in deep features, channels contain richer and diverse contextual information, which is mostly ignored by these methods. Therefore, we design a spatial and channel-based parallel scanning approach to compensate for the missing channel information in planar scanning, as shown in Fig. 3. We perform scan expansion and scan fusion on the $ch\times H$ plane. First, the input features are expanded on the $ch\times H$ plane and scanned from four directions to obtain the sequence $X_{1},X_{2},X_{3},X_{4}$. Then, in Channel-wise Selective Scanning, $B,C\in\mathbb{R}^{bs\times N\times L_{ch\times H}}$, and $\Delta\in\mathbb{R}^{bs\times d\times L_{ch\times H}}$ is calculated by Eq. 1-3 as an indication of the correlation between the feature channels. Finally, the sequences are merged and reconstructed into feature maps with the same size for images.
Image features are spatially selected by activating the mapping of each pixel in the spatial plane and the four directions of up, down, left and right to obtain local texture and structure information. In channel selection, the mapping relationship between different channels can be obtained to enhance the representation of similar channels in the implicit space in order to construct a wide range of global feature mappings and to enhance the scope of the sensory field. We visualized the range of effective receptive fields for different attentions, as shown in Fig. 1. As can be seen in Fig. 1 (d), the novel state space spatial channel attention module is able to maintain a larger and deeper receptive field and global feature mapping capability with a reduced number of parameters and linear complexity. This is consistent with our motivation to expect that enhancing the effective receptive field can better capture hand-object information associations from image features.
Figure 3: S3C module. In the S3C module, image features are fed into Spatial-wise Selective Scanning and Channel-wise Selective Scanning respectively. The correlation of spatial and channel dimensions in the hidden space is computed by selective state space modeling respectively. Eventually, they are merged to activate hand region features to obtain a feature map.
3.4 Regressor
@@
In Regressor, we input the hand attention features obtained from S3C into it. First, a 2D heat map of each joint $H$ is output through a simple Conv cascade. Then, a cascade of hand attention features and 2D heatmap $HM$ is input into four residual blocks he2015deep. Finally, the outputs of the residual blocks are vectorized into 2048-dimensional vectors and passed to the fully-connected layer for predicting the pose parameter $\theta\in\mathbb{R}^{48}$ and shape parameter $\mathfrak{\beta}\in\mathbb{R}^{10}$ of the MANO. At the same time, the joint regression matrices are multiplied by the original 3D mesh $\overline{\bar{T}}$ , and the forward kinematics are applied to obtain the final 3D hand joint coordinates $J\in\mathbb{R}^{21\times 3}$ and the 3D hand mesh $V\in\mathbb{R}^{778\times 3}$.
@@
To train our proposed HandS3C hand mesh reconstruction model, we constrain the model by calculating L2 loss on MANO parameters, 2D hand keypoints, 3D hand keypoints and vertex coordinates.
$\mathrm{Loss}=\alpha_{\theta}\|\theta^{\prime}-\theta\|+\alpha_{\beta}\|\beta^{\prime}-\beta\|+\alpha_{2d}\|J_{2d}^{\prime}-J_{2d}\|+\alpha_{3d}\|J_{3d}^{\prime}-J_{3d}\|+\alpha_{V}\|V^{\prime}-V\|$ (4)
@@
Table 1: COMPARISON WITH STATE-OF-THE-ART METHODS ON FREIHAND.
Method Params↓ PA-MPJPE↓ PA-MPVPE↓ F@5↑ F@15↑
Hasson et al. Hasson2019LearningJR 33.5M 13.3 13.3 0.429 0.907
ExPose Choutas2020MonocularEB 11.8 12.2 0.484 0.918
Kulon et al. kulon2020weaklysupervised 519.0M 8.4 8.6 0.614 0.966
I2L-MeshNet moon2020i2lmeshnet 136.8M 7.4 7.6 0.681 0.973
Pose2Mesh choi2021pose2mesh 75.0M 7.4 7.6 0.683 0.973
Tang et al. 9710830 156.5M 6.7 6.7 0.724 0.981
Junhyeong et al. cho2022crossattention 153.0M 6.5 0.982
HandS3C (Ours) 31.8M 7.6 7.0 0.690 0.982
Table 2: COMPARISON WITH STATE-OF-THE-ART METHODS ON DEXYCB.
Method Params↓ MPJPE↓ PA-MPJPE↓
Spurr et al. spurr2020weakly 17.34 6.83
METRO lin2021endtoend 229.5M 15.24 6.99
Liu et al. liu2021semisupervised 34.9M 15.28 6.58
HandOccNet park2022handoccnet 39.3M 14.04 5.80
H2ONet 10204388 - 14.00 5.70
Feng et al. Feng20243DHR - 13.20 5.60
HandGCAT Wang_2023 75.8M 13.76 5.60
HandS3C (Ours) 31.8M 12.81 5.52
Table 3: COMPARISON WITH STATE-OF-THE-ART METHODS ON HO3D.
I2L-MeshNet moon2020i2lmeshnet 136.8M 11.2 0.775 13.9 0.722 0.409 0.932
Hasson et al. hasson2020leveraging 13.2M 11.0 0.780 11.2 0.777 0.464 0.939
Hampali et al. hampali2020honnotate 10.7 0.788 10.6 0.790 0.506 0.942
METRO lin2021endtoend 229.5M 10.4 0.792 11.1 0.779 0.484 0.946
Liu et al. liu2021semisupervised 34.9M 9.9 0.803 9.5 0.810 0.528 0.956
I2UV-HandNet chen2021i2uvhandnet 9.9 0.804 10.1 0.799 0.500 0.943
ArtiBoost li2022artiboost 25.3M 11.4 0.773 10.9 0.782 0.488 0.944
Keypoint Hampali2021KeypointTS 56.5M 10.8 0.786
HandS3C (Ours) 31.8M 9.5 0.808 9.5 0.810 0.518 0.955
4 Experiments
4.1 Implementation Details
All implementations were performed with PyTorch. We used the Adam optimizer kingma2017adam, and each model was trained by annealing every 10th time starting from an initial learning rate of 10-4 to train the model with batch size of 32. Both our training and testing were performed on a server with an GPU RTX A5000 (24GB). All other details will be provided in our code.
4.2 Datasets and Evaluation Metrics
FREIHAND.
This dataset zimmermann2019freihand presents the first massively diverse hand dataset containing 32,560 training samples with the possibility of replacing four backgrounds. Also, there are 3960 test samples under occlusion and poor lighting.
DEXYCB.
This dataset chao2021dexycb is a large-scale dataset for object grasping and contains 582K image frames grasped on 20 YCB objects. We use the official “S0” segmentation to evaluate the right hand pose.
HO3D.
This dataset hampali2020honnotate is a hand-object interaction dataset with severe occlusion and contains 66,034 training samples and 11,524 test samples. The dataset provides RGB images with MANO-based hand joints and meshes and camera parameters. The test set provides only the annotations of the RGB images and the hand bounding box, and the test results can be evaluated through an online submission system.
Evaluation metrics.
For FreiHand, we report PA-MPJPE and PA-MPVPE in $mm$ as well as F-scores. For HO3D, we report PA-MPJPE/MPVPE, AUC and F-scores returned from the official evaluation server. For DexYCB, we report MPJPE and PA-MPJPE in $mm$.
Table 4: ABLATION RESULTS FOR THE FREIHAND.
Method PA-MPJPE↓ PA-MPVPE↓ F@5↑ F@15↑
@@ 4.2 Datasets and Evaluation Metrics
Baseline+S3C 7.6 7.0 0.690 0.982
Table 5: ABLATION RESULTS FOR THE DEXYCB.
Method MPJPE↓ PA-MPJPE↓
@@
Baseline+S3C 12.81 5.52
Table 6: ABLATION RESULTS FOR THE HO3D.
Method PA-MPJPE↓ PA-MPVPE↓ F@5↑ F@15↑
Baseline 9.8 9.7 0.516 0.954
Baseline+Self-Attention (4.6M) 10.1 10.1 0.502 0.947
Baseline+SSA (0.57M) 9.8 9.8 0.521 0.952
@@
Baseline+S3C (0.58M) 9.5 9.5 0.518 0.955
4.3 Comparisons with the State-of-the-art Methods

### v4

@@ Abstract
Reconstructing the hand mesh from one single RGB image is a challenging task because hands are often occluded by other objects. Most previous works attempt to explore more additional information and adopt attention mechanisms for improving 3D reconstruction performance, while it would increase computational complexity simultaneously. To achieve a performance preserving architecture with high computational efficiency, in this work, we propose a simple but effective 3D hand mesh reconstruction network (i.e., HandS3C), which is the first time to incorporate state space model into the task of hand mesh reconstruction. In the network, we design a novel state-space spatial-channel attention module that extends the effective receptive field, extracts hand features in the spatial dimension, and enhances regional features of hands in the channel dimension. This helps to reconstruct a complete and detailed hand mesh. Extensive experiments conducted on well-known datasets facing heavy occlusions (such as FREIHAND, DEXYCB, and HO3D) demonstrate that our proposed HandS3C achieves state-of-the-art performance while maintaining minimal parameters. Code can be available in https://github.com/JiaoZixun/HandS3C.
Index Terms:
3D Hand Mesh Reconstruction, Deep Learning, Effective Receptive Field, Human-computer Interaction, State Space Model.
I Introduction
@@
The reconstruction of 3D hand mesh from one single RGB image has a wide range of applications in many fields, such as VR/AR, robotics, and human-computer interaction and has attracted much attention due to its low cost and computational friendliness. Influenced by the rapid development of deep learning techniques, many excellent works [1, 2, 3, 4, 7, 5, 6] have emerged for this task. These works focus on important issues related to hand-object occlusion and the loss of local information. Despite significant advancements, most methods still struggle to reconstruct hand poses in some scenarios with substantial occlusion or when the prior knowledge is not sufficient.
@@
The previous methods can be broadly categorized into two categories. One is to incorporate additional prior knowledge [7, 8, 9, 10]. To cite few, Wang et al. [7] added keypoint coordinates as strong prior information to constrain hand pose. Zimmermann et al. [8] obtained hand region features via a hand segmentation network. Xie et al. [9] combined physiological aspects of hand motion to generate kinematic-compliant hand postures. These methods can enhance pose reconstruction performance, but lack essential prior information in practical applications (e.g., VR/AR games, robotics). Another class of approaches is to learn potential associations of image features through deeper architectures, but inevitably there is a scarcity of real annotation data. Some researchers [5, 11, 12] use Transformer architecture to obtain global information and expect to learn hand feature mapping from occluded parts. Other researchers [13] aggregated a large number of real datasets containing annotations for exploring associations between image features. However, these studies still have some limitations. First, the attention mechanism leads to excessive computational complexity. Second, an overload of global information can interfere with the effective receptive field [14].
@@
Figure 1: Overall architecture of HandS3C including Backbone, S3C and Regressor. In Backbone, ResNet-50 based encoder-decoder is employed to extract image features and fuses multiscale features. In S3C, image features are fed into Spatial-wise Selective Scanning and Channel-wise Selective Scanning modules, respectively. In Regressor, heat map features are obtained using convolution operation . Further MANO parameters are obtained from the FC layer to generate the hand mesh.
Based on the above observations, we propose a simple but effective 3D hand mesh reconstruction network called as HandS3C. This framework incorporates the state space model [15, 16, 17, 18] into the field of hand mesh reconstruction for the first time. At the same time, combining the idea of spatial-channel feature fusion [19], State Space Spatial Channel (S3C) attention module is constructed to emphasize the contribution of channel-dimensional features. This not only increases the range of effective receptive fields in the network, but also effectively avoids the square-step computational complexity of traditional attention. Specifically, our main contributions are summarized as follows:
  • We design a 3D hand mesh reconstruction network called as HandS3C, which introduces the state space model to the task of hand mesh reconstruction for the first time. This network can effectively improve hand reconstruction performance without the need for additional prior knowledge.
  • We propose a spatial and channel-wise parallel scanning approach to compensate for the missing channel information in planar scanning. The state-space spatial-channel attention module is constructed, which can enhance the effective receptive field range while maintaining small number of parameters.
  • Our method achieves state-of-the-art performance on three datasets, FREIHAND, DEXYCB and HO3D, with smaller number of parameters.
II Methodology
II-A Overview
@@
The overall network architecture is shown in Fig. 1 and consists of Backbone, S3C and Regressor components. Given a hand image $I\in\mathbb{R}^{256\times 256\times 3}$, the appearance information is firstly extracted by ResNet-50 based encoder-decoder and the multi-scale features are fused layer by layer to get the original feature map $F\in\mathbb{R}^{256\times 32\times 32}$. Then, it is fed into the S3C module and scanned in both spatial and channel dimensions to compute a 2D feature map and fused to obtain enhanced hand features $H\in\mathbb{R}^{256\times 32\times 32}$. Finally, the hand features $H$ are passed to the Regressor to calculate the keypoint heat map $HM\in\mathbb{R}^{21\times 32\times 32}$ to predict the keypoint coordinates. For the final 3D hand mesh, we use the most popular MANO parametric hand [20]. It is an articulated hand network generated from differentiable functions that takes as input the parameters that control the shape $\beta\in\mathbb{R}^{10}$ and pose $\theta\in\mathbb{R}^{48}$ of the hand.
II-B Backbone
@@
In the original feature extraction part, we use ResNet-50 to extract image information and fuse multi-scale features layer by layer in an U-shape encoder-decoder structure. Specifically, when the input image size is $256\times 256\times 3$, the size of the first layer of features after ResNet-50 becomes $128\times 128\times 128$. In the subsequent layers, each layer of image features $F_{i}\in\frac{H_{i}}{2^{i}}\times\frac{W_{i}}{2^{i}}\times C_{i},i\in(1,2,3,4),C_{i}\in(256,512,1024,2048)$ is obtained by downsampling. Then, upsampling is performed and the image features $F_{i}$ are incorporated layer by layer to get the last layer of feature extraction with size of $256\times 64\times 64$. Finally, the feature mapping map $F\in\mathbb{R}^{256\times 32\times 32}$ is obtained by fusing the feature information through average pooling.
II-C State Space Spatial Channel (S3C) Attention Module
Spatial-wise Selective Scanning
Unlike textual information, image information contains 2D spatial information such as local texture and global structure, and it is difficult to obtain the contextual relationship of the local sense field of an image by scanning with a simple S6 module. The 1-D convolution is extended to 2-D convolution in S4ND [21], and this approach leads to the loss of the ability to dynamically update weights based on context. In VMamba [15], rows and columns are unfolded and scanned along in four different directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right, by means of a cross-scanning module. This allows any pixel to contain contextual information from all four directions and can enhance the size of the sensory field in the $H\times W$ plane. Sequences in each of the four directions are computed in the Spatial-wise Selective Scanning to obtain $B,C\in\mathbb{R}^{bs\times N\times L_{H\times W}}$, and $\Delta\in\mathbb{R}^{bs\times d\times L_{H\times W}}$, which in turn computes the feature map. Finally, the four sequences are merged to obtain a new sequence with the same size as the input sequence.
Channel-wise Selective Scanning
Although spatial features such as local texture and global structure intrinsic to 2D image features are already available through Liu et al [15]. However, in deep features, channels contain richer and diverse contextual information, which is mostly ignored by these methods. Therefore, we design a spatial and channel-based parallel scanning approach to compensate for the missing channel information in planar scanning, as shown in Fig. 1. We perform scan expansion and scan fusion on the $ch\times H$ plane. First, the input features are expanded on the $ch\times H$ plane and scanned from four directions to obtain the sequence $X_{1},X_{2},X_{3},X_{4}$. Then, in Channel-wise Selective Scanning, $B,C\in\mathbb{R}^{bs\times N\times L_{ch\times H}}$, and $\Delta\in\mathbb{R}^{bs\times d\times L_{ch\times H}}$ is calculated by Eq. 1-3 as an indication of the correlation between the feature channels. Finally, the sequences are merged and reconstructed into feature maps with the same size for images.
Image features are spatially selected by activating the mapping of each pixel in the four directions to obtain local texture and structure information. In channel selection, the mapping relationship between different channels can be obtained to enhance the representation of similar channels in the implicit space in order to construct a wide range of global feature mappings and to enhance the scope of the sensory field. The novel state space spatial channel attention module is able to maintain a larger and deeper receptive field and global feature mapping capability with a reduced number of parameters and linear complexity. This is consistent with our motivation to expect that enhancing the effective receptive field can better capture hand-object information associations from image features.
@@ II-C State Space Spatial Channel (S3C) Attention Module
$h_{i}=$ $\overline{A}h_{i-1}+\overline{B}x_{i},y_{i}=Ch_{i}+Dx_{i},$ (2)
$\overline{A}=$ $e^{\Delta A},\overline{B}=(e^{\Delta A}-I)A^{-1}B,\overline{C}=C$
@@
II-D Regressor
@@
In Regressor, we input the hand attention features obtained from S3C into it. First, the 2D joint heat map $HM$ is computed using a convolution operation. Then, a cascade of hand attention features and a 2D heatmap $HM$ is input into four residual blocks. Finally, the pose parameter $\theta\in\mathbb{R}^{48}$ and the shape parameter $\mathfrak{\beta}\in\mathbb{R}^{10}$ of the MANO are predicted. Detailed 3D hand joint coordinates $J\in\mathbb{R}^{21\times 3}$ and 3D hand mesh $V\in\mathbb{R}^{778\times 3}$ are obtained from the MANO hand model generated.
@@
To train our proposed HandS3C model for hand mesh reconstruction, we constrain the model by calculating L2 loss.
$\mathrm{Loss}=\alpha_{\theta}\|\theta^{\prime}-$ $\theta\|+\alpha_{\beta}\|\beta^{\prime}-\beta\|+\alpha_{2d}\|J_{2d}^{\prime}-J_{2d}\|$ (4)
$+$ $\alpha_{3d}\|J_{3d}^{\prime}-J_{3d}\|+\alpha_{V}\|V^{\prime}-V\|$
@@
TABLE I: COMPARISON WITH STATE-OF-THE-ART METHODS ON FREIHAND.
Method Params↓ PA- PA- F@5↑ F@15↑
MPJPE↓ MPVPE↓
Hasson et al. [1] 33.5M 13.3 13.3 0.429 0.907
ExPose [3] 11.8 12.2 0.484 0.918
Kulon et al. [22] 519.0M 8.4 8.6 0.614 0.966
I2L-MeshNet [23] 136.8M 7.4 7.6 0.681 0.973
Pose2Mesh [24] 75.0M 7.4 7.6 0.683 0.973
Tang et al. [25] 156.5M 6.7 6.7 0.724 0.981
Junhyeong et al. [26] 153.0M 6.5 0.982
AHRNET [27] 7.2 0.978
@@
HandS3C (Ours) 31.8M 7.0 6.3 0.727 0.989
TABLE II: COMPARISON WITH STATE-OF-THE-ART METHODS ON DEXYCB.
Method Params↓ MPJPE↓ PA-MPJPE↓
Spurr et al. [28] 17.34 6.83
METRO [29] 229.5M 15.24 6.99
Liu et al. [30] 34.9M 15.28 6.58
HandOccNet [5] 39.3M 14.04 5.80
H2ONet [31] - 14.00 5.70
Feng et al. [32] - 13.20 5.60
HandGCAT [7] 75.8M 13.76 5.60
MS-MANO [9] 12.92
@@
HandS3C (Ours) 31.8M 12.81 5.52
TABLE III: COMPARISON WITH STATE-OF-THE-ART METHODS ON HO3D.
I2L-MeshNet [23] 136.8M 11.2 0.775 13.9 0.722 0.409 0.932
Hasson et al. [38] 13.2M 11.0 0.780 11.2 0.777 0.464 0.939
Hampali et al. [35] 10.7 0.788 10.6 0.790 0.506 0.942
METRO [29] 229.5M 10.4 0.792 11.1 0.779 0.484 0.946
Liu et al. [30] 34.9M 9.9 0.803 9.5 0.810 0.528 0.956
I2UV-HandNet [36] 9.9 0.804 10.1 0.799 0.500 0.943
ArtiBoost [37] 25.3M 11.4 0.773 10.9 0.782 0.488 0.944
Keypoint [4] 56.5M 10.8 0.786
Baseline 9.8 0.798 9.7 0.801 0.516 0.949
Baseline+Self-Attention(4.6M) 10.1 0.798 10.1 0.798 0.502 0.947
Baseline+SSA(0.57M) 9.8 0.803 9.8 0.804 0.521 0.952
@@
Baseline+S3C(0.58M) (Ours) 31.8M 9.5 0.808 9.5 0.810 0.518 0.957
III Experiments
III-A Implementation Details
We train on three publicly available datasets, FREIHAND[33], DEXYCB[34], and HO3D[35], and the model was trained by annealing every 10th time starting from an initial learning rate of 10-4 to train the model with batch size of 32. Our training and testing were performed on a server equipped with a GPU RTX A5000 (24GB). All other details will be provided in our code. In the FREIHAND, we report PA-MPJPE and PA-MPVPE and F-scores in $mm$. For DEXYCB, we report MPJPE and PA-MPJPE in $mm$. For HO3D, we report PA-MPJPE/MPVPE, AUC and F scores returned from the official evaluation server.
III-B Comparisons with the State-of-the-art Methods